Solve (3x+y-z)p + (x+y-z)q = 2(z+y)
1
Expert's answer
2021-10-12T15:18:54-0400

"\frac{dx}{3x+yz}=\frac{dy}{x+yz}"

"(x+yz)dx=(3x+yz)dy"

"\frac{x^2}{2}+xyz-3xy-\frac{y^2z}{2}=C_1"

"\frac{dy}{x+yz}=\frac{dz}{2(z+y)}"

"(x+yz)dz=2(z+y)dy"

"xz+\frac{yz^2}{2}-2zy-y^2=C_2"

Answer:"\psi(\frac{x^2}{2}+xyz-3xy-\frac{y^2z}{2};xz+\frac{yz^2}{2}-2zy-y^2)" =0